OpenStudy (anonymous):

PLEASE HELP- ALGEBRA/PROBABILITY QUESTION: 5 white balls and k black balls are placed into a bin. Two of the balls are drawn at random. The probability that one of the drawn balls in white and the other is black is 10/21. Find the smallest possible value of k.

OpenStudy (uri):

@terenzreignz Is bored,He will help you.

OpenStudy (ybarrap):

solve - \[5k/((5+k(4+k) + 5k/((5+k)(5+k-1) = 10/21\] you will get k=2 or k=10 So smallest k is 2

OpenStudy (anonymous):

Thanks, I believe that is correct

OpenStudy (ybarrap):

The two terms of the sum are probability of getting white black or black white

OpenStudy (anonymous):

Cause 5k= choosing 1 white and 1 black, then I *think* you can find the total number of choosing two by doing 7c2.

OpenStudy (anonymous):

So 5k= 10, and 7 c 2 = 21

OpenStudy (ybarrap):

Equation should be \[5k/((5+k)(4+k)) + 5k/((5+k)(5+k-1)) = 10/21\]

OpenStudy (ybarrap):

\[(1/2)*5k/((5+k)(4+k))=10/21 \]

OpenStudy (ybarrap):

The rest is algebra, quadratic equation stuff

OpenStudy (ybarrap):

That last equation should have been: \[2∗5k/((5+k)(4+k))=10/21 \]

OpenStudy (ybarrap):

Because both P(wb) and P(bw) are equal

OpenStudy (ybarrap):

All the algebraic steps are here: Solve for k over the real numbers: (10 k)/((k+4) (k+5)) = 10/21 Multiply both sides by a polynomial to clear fractions. Cross multiply: 210 k = 10 (k+4) (k+5) Write the quadratic polynomial on the right hand side in standard form. Expand out terms of the right hand side: 210 k = 10 k^2+90 k+200 Move everything to the left hand side. Subtract 10 k^2+90 k+200 from both sides: -10 k^2+120 k-200 = 0 Factor the left hand side. The left hand side factors into a product with three terms: -10 (k-10) (k-2) = 0 Divide both sides by a constant to simplify the equation. Divide both sides by -10: (k-10) (k-2) = 0 Solve each term in the product separately. Split into two equations: k-10 = 0 or k-2 = 0 Look at the first equation: Solve for k. Add 10 to both sides: k = 10 or k-2 = 0 Look at the second equation: Solve for k. Add 2 to both sides: Answer: | | k = 10 or k = 2

OpenStudy (ybarrap):

Remember, you are drawing a ball without replacement. So the the second term (representing the second ball drawn) in either P(wb) or P(bw) will have one less ball in the denominator

OpenStudy (ybarrap):

Does this make sense?